OpenStudy (anonymous):

l:b::7:5 and area is 1260cm² find the perimeter of the rectangle

OpenStudy (anonymous):

@megangray32 @JayDS @BasedGod1122

OpenStudy (jtvatsim):

l:b :: 7:5 means that l = 7x and b = 5x for some unknown

OpenStudy (jtvatsim):

if you plug that into a rectangle, use the formula for area and set that equal to 1260, you should have your answer.

OpenStudy (anonymous):

35X = 1260CM²

OpenStudy (jtvatsim):

35x^2 watch for the squared. :)

OpenStudy (anonymous):

WHAT?

OpenStudy (jtvatsim):

(7x)(5x) = 35x^2

OpenStudy (jtvatsim):

do you see that? :)

OpenStudy (anonymous):

okay now waht

OpenStudy (jtvatsim):

first get rid of that pesky 35 by division

OpenStudy (anonymous):

how?

OpenStudy (jtvatsim):

you have 35x^2 = 1260 right? So divide both sides by 35.

OpenStudy (anonymous):

6?

OpenStudy (jtvatsim):

almost, not quite since 35 * 6 = 210 :)

OpenStudy (anonymous):

36?

OpenStudy (jtvatsim):

good! :)

OpenStudy (jtvatsim):

so now, x^2 = 36 any ideas?

OpenStudy (anonymous):

so is x 6?

OpenStudy (jtvatsim):

yahoo! you got it!

OpenStudy (jtvatsim):

but we aren't done yet... :)

OpenStudy (anonymous):

so l is 42 and b is 30?

OpenStudy (jtvatsim):

good, so the perimeter is?

OpenStudy (anonymous):

144?

OpenStudy (jtvatsim):

yep yep!

OpenStudy (jtvatsim):

nicely done

OpenStudy (anonymous):

what is the units for perimeter 144 cm feet or what

OpenStudy (jtvatsim):

well, the area is cm^2, so we have to be something in cm, and length is measured in cm.

OpenStudy (jtvatsim):

so 144 cm like you said.